OpenStudy (anonymous):

what is tan(pi/8) rounded to 4 decimal places what is tan(pi/8) rounded to 4 decimal places @Mathematics

OpenStudy (anonymous):

pls show work

OpenStudy (mathteacher1729):

\(\sqrt{2}-1\) Do they want you to use any special means to compute the square root of two?

OpenStudy (anonymous):

they just say find the value rounded to 4 decimal places

OpenStudy (eyust707):

plug it into your calculator. then round it

OpenStudy (mathteacher1729):

So you can use the tangent half-angle formula here... pi/8 is just one half of pi/4. Here is the diagram for the tangent half-angle formula: http://en.wikipedia.org/wiki/File:Weierstrass_substitution.png

OpenStudy (anonymous):

ok what about eyust method...can i actually do that

OpenStudy (eyust707):

it's up to your instructor

OpenStudy (eyust707):

bsed on the stuff we were doing earlier than they probably are just tryoing to teach you how to use your calc in this case

OpenStudy (anonymous):

ok what if i have sec(97) theres no button for sec on the calc...what do i do then

OpenStudy (eyust707):

well sec x = 1/cos x

OpenStudy (eyust707):

i could show you why if you would like

OpenStudy (eyust707):

but thats a property you should memorize

OpenStudy (anonymous):

yea pls go ahead

OpenStudy (eyust707):

honestly i cant quite remember.. it has to do with the properties of the unit circle.. your teacher may have derived it.. maybe mathteacher can show you?

OpenStudy (anonymous):

mathteacher are you there?

OpenStudy (mathteacher1729):

Yes I'm still here sorry.

OpenStudy (eyust707):

but for now just memorize these: http://www.mathacademy.com/pr/prime/articles/trig_ident/trig_ident_6.gif

OpenStudy (eyust707):

why does secx = 1/cosx?

OpenStudy (mathteacher1729):

eyust -- that is just a definition. It is defined to be that way.

OpenStudy (anonymous):

using that definition, mathteacher, how can i solve sec(97)

OpenStudy (eyust707):

sec(97) = 1/cos(97)

OpenStudy (mathteacher1729):

You wouldn't "solve" it, you would "evaluate" it. And since 97 is not a "nice" angle (pi/2, pi/3, or some combination of those), then we just use a calc, or some more advanced numerical approximations.

OpenStudy (anonymous):

so our final evaluation would just be 1/cos(97)?

OpenStudy (eyust707):

yep

OpenStudy (eyust707):

is that 97 in degrees or radians?

OpenStudy (anonymous):

degrees

OpenStudy (eyust707):

make sure your calc is set to degrees when you type it in

OpenStudy (anonymous):

o ok...when i type it in i get -8.205509...4 decimals would be -8.2055...thats it?

OpenStudy (anonymous):

mathteacher,is that it?